OpenStudy (alymorrison):

A sphere has a diameter of 14 units. What is the volume of the sphere in cubic units? If a cylinder has the same radius as the sphere and a height of 14 units, what is the volume of the cylinder? Use 3.14 for π.

jimthompson5910 (jim_thompson5910):

`A sphere has a diameter of 14 units` so the radius is 14/2 = 7 units r = 7

jimthompson5910 (jim_thompson5910):

Volume of a Sphere \[\Large V = \frac{4}{3}\pi*r^3\] where \(\Large \pi \approx 3.14\) and r = 7 in this case

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (alymorrison):

-The volume of the sphere is about 1,077.02 cubic units, and the volume of the cylinder is about 718.01 cubic units. -The volume of the sphere is about 1,436.03 cubic units, and the volume of the cylinder is about 2,154.04 cubic units. -The volume of the sphere is about 1,436.03 cubic units, and the volume of the cylinder is about 957.35 cubic units. -The volume of the sphere is about 1,077.02 cubic units, and the volume of the cylinder is about 1,615.53 cubic units.

OpenStudy (alymorrison):

those are my choices

jimthompson5910 (jim_thompson5910):

does the formula I posted look familiar? have you used it before?

OpenStudy (alymorrison):

no sir

jimthompson5910 (jim_thompson5910):

Replace every copy of 'r' with 7, then evaluate, to get \[\Large V = \frac{4}{3}\pi*r^3\] \[\Large V = \frac{4}{3}*3.14*(7)^3\] \[\Large V \approx 1,436.0267\] This volume is approximate. I used a calculator on the last step. I typed in `(4/3)*3.14*7^3`

jimthompson5910 (jim_thompson5910):

If you don't have a calculator, I recommend this one http://web2.0calc.com/

OpenStudy (alymorrison):

thank you

jimthompson5910 (jim_thompson5910):

for the cylinder, you'll use this formula \[\Large V = \pi*r^2*h\] where r = radius h = height